If the process can be assumed to follow a Gaussian distribution then 99.7% of the outputs of the process will lie between those two limits. That may be of benefit in quality control if it is a production process.
What is the probability of landing on heads twice?
The probability of landing on heads each time a fair coin is flipped, is 1/2.
Assuming that the question was supposed to be:
"What is the probability of landing on heads twice in a row?"
To calculate compound probabilities like this, we first have to work out the probability of landing on heads each time, and then multiply these two probabilities to get a compound probability.
1/2 x 1/2 = 1/4
So the probability of landing on heads twice in a row = 1/4 (for a fair coin)
How do you find the sample size when given the standard deviation and the mean with a sample value?
You cannot from the information provided.
What is the probability of drawing a marble that is not white?
It depends on how many marbles of each colour you have....
Does the probability of the type 1 error increase as the sample size increases?
No, the estimates should become more robust and the power of the test should, therefore, increase.
How many different ways can you arrange the word history?
The h can be in any of seven place, the i in any of the remaining six etc etc so 7 x 6 x 5 x 4 x 3 x 2 = 5040 different ways.
How many different ways can you arrange 7 letter words?
The best way to think of this is how many options there are for each letter. There are 7 possibilities for the first letter. Then there are 6 possibilities (out of the remaining letters) for the second latter. Then 5 possibilities for the next, and so on. The total number of possibilities will be 7x6x5x4x3x2x1, which equals 5,040. So there are 5,040 different ways to arrange 7 letter words.
What is the probability that in a room of 8 people 2 have the same birthday?
P = 0.072314699... ≈ 7.23%
SOLUTION
Let's identify the 8 people by integers 1, 2,..., 8, and their birthday by x1, x2,..., x8.
To include the possibility of xi = Feb. 29 (of a leap year), we will consider 4(365) +
1 = 1451 days, so the probability of a person having Feb. 29 as his birthday is
P(xi = Feb. 29) = 1/1461. The probability of a person having his birthday on any
given day except Feb. 29 is P(xi ≠ Feb. 29) = 4/1461.
Let's first calculate the probability of person 1 and 2 having the same birthday, x1 = x2, and persons 3, 4,..., 8, all have different birthdays x3,...,x8.
Now we go.
We take person 1 and have two possibilities:
A).- x1 = Feb. 29 B).- x1 ≠ Feb. 29
P(x1 = Feb. 29) = 1/1461; P(x ≠ Feb. 29) = 1 - 1/1461
We will construct two branches. One for each possibility.
Branch A: given x1 = Feb. 29, the probability of person 2 having the same birthday
is P(x2 = x1) = 1 - 1/1461. The probability of person 3 not having the same birthday is P(x3 ≠ x1 = x2) = 1 - 1/1461. The probability of person 4 not having either two birthdays is P(x4 ≠ x1, x4 ≠ x3) = 1 - 5/1461. The probability of person 5 not sharing x1 nor x3 nor x4 is P(x5 ≠ x4, x5 ≠ x3, x5 ≠ x2) = 1 - 9/1461. And so a series of terms develops:
P(xi ≠ ... ) = 1 - (4i - 3)/1461 from i = 1, to i = n-2
where n is the number of people in the room (In this particular question for 8
people, n = 8).
The branch we have constructed for possibility A is the product of the individual
conditional probabilities;
A: (1/1461)2 π1n-2[1-(4i-3)/1461]
where π1n-2 is the product operator "pi" of "i" terms from i =1, to i = n-2.
Branch B: Given x1 ≠Feb. 29, the probability of person 2 having the same birthday
is P(x2 = x2) = 4/1461. The probability of person 3 not having the same birthday
is P(x3 ≠ x1) = 1 - 4/1461. The probability of person 4 not sharing birthdays with 2 and 3 is P(x4 ≠ x3, x4 ≠ x2) = 1 - 8/1451. The probability of person 5 not
sharing x2 nor x3 nor x4 is P(x5 ≠ ...) = 1 - 12/1461. And so a series of terms
develops:
P(xi ≠ ... ) = 1- (4i -4)/1461 from i = 2, to i = n-1
where n is the number of people in the room (In this particular question for 8 people, n = 8).
Branch for possibility B: (1-1/1461)(4/1461) π2n-1[1-(4i-4)/1461]
The sum of these two branch expressions give the probability of persons 1 and 2
sharing a birthday in a group of n persons:
P(x1=x2)=(1/1451)2π1n-2[1-(4i-3)/1461]+(1-1/1461)(4/1461)π2n-1[1-(4i-4)/1461]
Now, the probability of any two persons, and only those two persons, to share a birthday in a group of n persons, is the previous probability, times the number of different combinations of two persons, (i,j), in a group of n persons, nC2
(where, nC2 = n!/[2!(n-2)!] )
The final expression is:
--------------------------------------------------------------------------------------------
P = nC2{(1/1451)2π1n-2[1-(4i-3)/1461]+(1-1/1461)(4/1461)π2n-1[1-(4i-4)/1461]}
--------------------------------------------------------------------------------------------
For the case of n = 8, the result is, P = 0.072314699... ≈ 7.23%
On an unweighted die, regardless of how many successes you've had, your chance of rolling a five will be one in six.
The chance of you rolling a die 16 times and getting a five each time would be 1/616, or 0.00000000000035447042.
The chance of rolling a die 16 times and getting the same number (regardless of what that number is) each time, would be 1/615, or 0.00000000000212682249
Regardless of how slim that chance is though, your chance on the next roll will still be 1/6.
However, if this is meant as a "real world" question, then your chances of rolling the same number so many times in a row is so low that at that point, your odds would be much higher of there being something odd with the die, or with your experiment. At that point, it would be sensible to say that the odds are very good of rolling another 5, regardless of the math, as there seems to be another factor affecting your outcome.
What is the theoretical probability of a six sided die when it is rolled 24 times?
well its all 50% no matter what
What is the probability of rolling a 4 or an odd number on a die?
The probability of rolling a 4 is 1/6. (There are 6 numbers one of which is "4")
The probability of rolling an odd number is 3/6 = 1/2 (There are 6 numbers 3 of which are odd)
Thus, in one roll of the die, the odds of rolling a 4 oran odd number = 1/6 + 3/6 = 4/6 = 2/3.
What is the probability of spinning a prime number?
The answer depends on how many numbers on the spinner, and their distribution. Without that information, it is not possible to answer the question.
What is the probability of picking a red ace out of a standard deck of cards?
The probability of pick a red ace out of a standard deck of cards would be 1/26 because there are two read aces, the ace of diamonds and the ace of hearts.
What is the probability of picking an m from the word mathematics?
If picked at random, the answer is 2/11.
What would be the theoretical probability of rolling a die 50 times and rolling a 2?
The probability of rolling at least one 2 in fifty rolls of a standard die is 1 - (5/6) 50, or about 0.99989012.
This calculation starts by looking at the probability of not rolling a 2, which is 5/6. To repeat that 50 times in a row, you simply raise that to the 50th power, getting 0.000109885. Then you subtract the result from 1 to get the probability of not succeeding in not rolling a 2 in fifty tries.
Expressed in normal "odds" notation, this is about (100000 - 11) in 100000, or about 99989 in 100000.
What is the probability of choosing a marble that is not green in problem 2?
Since you haven't seen fit to tell us problem 2, we'd only be guessing.
How many red cards is there in a deck of cars?
yguywegoyu2df2ud24dig2cegdc2eougo3uyrtey3vdgegweccndbcagscdwggdwqgewdgwekgcwlhjfvvefggfgvwfgvwlgevkehwvhgwehefwvh;fwvehvfvhu;fvhfebhjmddhvbhvbhffhd8sdbefhu3v2y3vfbhvgfcsvgabHnjmkl